OpenStudy (anonymous):

if sin theta equals -5/13 pi is less than theta less than 3 pi/2 find cos theta over 2 This is just all over the place

OpenStudy (anonymous):

If \(\sin \theta = -\frac{ 5 }{ 13 }, \pi<\theta<\frac{ 3\pi }{ 2 }\), find \(\cos \frac{ \theta }{ 2 }\). this?

OpenStudy (anonymous):

haha yes!

OpenStudy (anonymous):

you need to use the half angle identity \[\cos \theta = \pm \sqrt{\frac{ 1+\cos \theta }{ 2}}\] which means you need to find cos Θ.

OpenStudy (anonymous):

Where are you suppose to use the sin theta part then?

OpenStudy (anonymous):

Θ is in the 3rd quadrant so cos Θ is negative, and you can use the pythagorean theorem/triples to find it

OpenStudy (anonymous):

Right so its 5 12 13 triangle what would you use for the cos theta part?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so cos Θ would be -12/13

OpenStudy (anonymous):

make sense?

OpenStudy (anonymous):

So then you would just solve after?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

Alright lets see how this works working it out then I might show you the work to see if its correct lol

OpenStudy (anonymous):

@peachpi Like this?

OpenStudy (anonymous):

I can't really see if the 12/13 is negative. If it is then you're right and you just have to simplify it

OpenStudy (anonymous):

Alright yeah its negative sorry its hard to tell but we got sqrt of 1 over 26?

OpenStudy (anonymous):

that's what I got. Then since Θ is in the third quadrant, Θ/2 is in the second where cos is negative so got with -√(1/26)

OpenStudy (anonymous):

or -(√26)/26 if you have to rationalize

OpenStudy (anonymous):

but I don't think either of them are right becauase the answer has to be greater then pi and less then 3 pi over 2?

OpenStudy (anonymous):

that's a restriction on Θ. That's how we know Θ is in the 3rd quadrant.

OpenStudy (anonymous):

the question is asking for cos (Θ/2), not Θ itself

OpenStudy (anonymous):

The answer is simple, A true saiyan sprinkles when he tinkles.

OpenStudy (anonymous):

Wait so whats the answer then? im so confused lmaoo

OpenStudy (anonymous):

I just said it

OpenStudy (anonymous):

cos (Θ/2) = -√(1/26)

OpenStudy (anonymous):

Is there a way to check the answer itself?

OpenStudy (anonymous):

try wolframalpha.com not sure how you'd have to word it though

OpenStudy (anonymous):

But are you sure your right? with the answer and everything?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Alright I have an exam coming up which is why I am asking these questins XD haha thank you!

OpenStudy (anonymous):

you're welcome